# Differentiable Annealed Importance Sampling Minimizes The Jensen-Shannon Divergence Between Initial and Target Distribution

### v1

@@ Abstract
Differentiable annealed importance sampling (DAIS), proposed by Geffner & Domke (2021) and Zhang et al. (2021), allows optimizing, among others, over the initial distribution of AIS. In this paper, we show that, in the limit of many transitions, DAIS minimizes the symmetrized KL divergence (Jensen-Shannon divergence) between the initial and target distribution. Thus, DAIS can be seen as a form of variational inference (VI) in that its initial distribution is a parametric fit to an intractable target distribution. We empirically evaluate the usefulness of the initial distribution as a variational distribution on synthetic and real-world data, observing that it often provides more accurate uncertainty estimates than standard VI (optimizing the reverse KL divergence), importance weighted VI, and Markovian score climbing (optimizing the forward KL divergence).
@@ 1 Introduction
We show theoretically that, in the limit of many transitions, DAIS minimizes the symmetrized Kullback-Leibler (KL) divergence (Kullback & Leibler, 1951), also known as Jensen-Shannon divergence, between its initial and target distribution. The Jensen-Shannon divergence is the sum of the reverse and the forward KL divergence. While the reverse KL divergence, used in VI, is known to be mode-seeking, the forward KL divergence, used in Markovian Score Climbing (MSC) (Naesseth et al., 2020), is typically associated with a mass-covering behavior (Murphy, 2023). The Jensen-Shannon divergence averages between both divergences.
@@
Starting from our theoretical contribution—showing that DAIS minimizes the Jensen-Shannon divergence between $q_{0}$ and ${f\,/\,Z}$—we provide an empirical analysis of $\text{DAIS}_{0}$ for approximate Bayesian inference. We compare $\text{DAIS}_{0}$, DAIS, VI, importance weighted VI (IWVI), and MSC. We find that $\text{DAIS}_{0}$ often provides more accurate uncertainty estimates than VI, IWVI, and MSC on Gaussian process regression and real-world datasets for logistic regression.
@@ 2 Related Work
In the following, we discuss related work on variational distributions that are augmented by MCMC transitions, the forward and reverse KL divergence, and DAIS. We cover related work on methods to estimate normalization constants in Section 3. Related work on bounds with respect to the Jensen-Shannon divergence is given in Section 4.
@@
Black box VI (Ranganath et al., 2014) is typically associated with the reverse KL divergence (between the variational distribution and the real posterior distribution). However, also various other divergences have been investigated to obtain an approximation of the posterior distribution (Hernandez-Lobato et al., 2016; Li & Turner, 2016; Dieng et al., 2017; Wan et al., 2020). Most related to $\text{DAIS}_{0}$, Ruiz & Titsias (2019) refine the variational distribution by running MCMC transitions. They minimize a contrastive divergence which, in the limit, converges to the Jensen-Shannon divergence. Markovian score climbing (MSC) (Naesseth et al., 2020), that we also compare to in Section 5, optimizes the forward KL divergence using unbiased stochastic gradients. MSC samples a Markov chain and uses the samples to follow the score of the variational distribution. The Markov kernel for the MCMC dynamics is based on importance sampling.
@@ 4 Analyzing the Initial Distribution of DAIS
In Theorem 4.1 of this section we present our main theoretical result that DAIS minimizes the Jensen-Shannon divergence between its initial distribution and its target distribution. Then, we discuss compact representations in terms of their sampling complexity. Finally, we frame finding the initial distribution of DAIS as a form of VI.
4.1 DAIS Minimizes the Jensen-Shannon Divergence
Theorem 4.1 below presents our main theoretical contribution showing that DAIS minimizes the Jensen-Shannon divergence between its initial and target distribution. We see Theorem 4.1 as starting point to motivate $\text{DAIS}_{0}$ and the analysis of the shape of the fitted initial distribution $q_{0}$. The result also holds true for AIS, but it is most relevant for DAIS, which learns parameters of its initial distribution. One can also show Theorem 4.1 by combining results of Grosse et al. (2013) and Crooks (2007) from the physics literature. Brekelmans et al. (2022) state a related bound on the difference between an upper and a lower bound on the evidence (whereas Theorem 4.1 makes an asymptotic statement about the gap between lower bound and evidence).
@@ 4.1 DAIS Minimizes the Jensen-Shannon Divergence
Let $\operatorname{support}(q_{0})\supset\operatorname{support}(f)$. We assume that the transitions between consecutive annealing distributions are perfect and that $\beta_{k}$ are equally spaced, i.e., $\beta_{k}=k/K$. Then, for large $K$ and $N=1$, the gap $\Delta_{\mathrm{AIS}}^{1,K}\coloneqq\log Z-\mathrm{ELBO}_{\mathrm{AIS}}^{1,K}(f_{\mathrm{AIS}},q_{\mathrm{AIS}})$ is a divergence between the initial distribution $q_{0}$ and the target distribution $f/Z$. Up to corrections of ${\mathcal{O}}(1\,/\,K^{3})$, the divergence is proportional to the Jensen-Shannon divergence,
@@
where $D_{\text{JS}}$ denotes the Jensen-Shannon divergence. This shows that the bound for $N>1$ can be upper-bounded by the Jensen-Shannon divergence but does not give additional insights on, e.g., symmetry. For $N=1$ the inequality is an equality. See Appendix A for a discussion in greater detail.
@@
In contrast, the forward KL divergence $D_{\mathrm{KL}}({f\,/\,Z}\,\|\,q)$ is mass-covering since the expectation is taken over $f/Z$ and regions in ${\bm{z}}$-space where $q({\bm{z}})<f({\bm{z}})\,/\,Z$ are weighted higher than the regions where $q({\bm{z}})>f({\bm{z}})\,/\,Z$. While the forward KL divergence is less likely to underestimate posterior variances, it often performs poorly in moderate to high dimensions when posterior correlations increase (Dhaka et al., 2021). Section 5 provides empirical evidence that $\text{DAIS}_{0}$ (implicitly minimizing the Jensen-Shannon divergence) indeed outperforms MSC (minimizing the forward KL divergence).
@@ 4.3 The Initial Distribution of DAIS for Inference
As highlighted in Section 4.2, we do not expect that the initial distribution of DAIS ($\text{DAIS}_{0}$) generally outperforms the DAIS method. However, we highlight that $\text{DAIS}_{0}$ provides various desirable properties due to its compact representation. Additionally, $\text{DAIS}_{0}$ is much more computationally efficient at inference time compared to DAIS. In comparison to VI, $\text{DAIS}_{0}$ is expected be less mode-seeking (Section 4.1), and we pose that the Jensen-Shannon divergence is less susceptible to problems optimizing the forward KL divergence. Section 5 provides experimental evidence.
@@ 5.1 Bimodal Target Distribution
In Figure 4 we can clearly see that VI and $\text{DAIS}_{0}$ with $K=2$ transitions lead to a mode-seeking distribution. For an increasing $K$ (i.e., $K\in\{4,16\}$), we find that the variational distribution of $\text{DAIS}_{0}$ becomes more mass-covering and increasingly similar to the solution that (numerically) minimizes the Jensen-Shannon divergence (regularly dashed line). This is consistent with Theorem 4.1. We find that we can unambiguously classify almost all of the learned variational distributions into being mode-seeking (“s”) or mass-covering (“c”) by calculating the distance of the mean of the variational distribution to both modes of the target distribution, and to their mid point (Appendix B provides a histogram of these distances and more details on the classification). Figure 3 shows corresponding classifications as labels “s” and “c” for each combination $(N,K,d)$. We label the few cases where the classification is not perfectly unambiguous with “u” which is short for for “undecidable”.
Comparing (IW)VI and $\text{DAIS}_{0}\ (16)$ in Figure 3, we find that $\text{DAIS}_{0}\ (16)$ typically covers both modes of the target distribution for larger dimensions (across all $N$). We attribute this finding to the Jensen-Shannon divergence that is implicitly minimized by $\text{DAIS}_{0}$. MSC does not converge to reasonable values (“-”) for $N=1$. With a single chain $\text{MSC}_{1\text{c}}$ outperforms (IW)VI over all $N$ but seems to be less mass-covering than $\text{DAIS}_{0}$. With $8$ parallel chains $\text{MSC}_{8\text{c}}$ sometimes outperforms the other methods in finding a mass-covering distribution. However, we find that the variational distributions of $\text{MSC}_{8\text{c}}$ are often not centered on the mid-point between the two modes which we attribute to the high variance of the method (see Appendix B). This is in accordance with Kim et al. (2022) who report high variance for MSC and Dhaka et al. (2021) who note that the forward KL divergence is difficult to optimize for large dimensions.
@@ 5.2 Gaussian Processes Regression
Table 1 shows mean absolute errors (MAEs) for a larger set of GP models. We compare models with compact distributions (IWVI, $\text{DAIS}_{0}$, $\text{MSC}_{8\text{c}}$) and sampling based approaches (DAIS using $10^{5}$ and $\text{IWVI}_{\text{SIR}}$ using $10^{3}\times 10^{3}$ samples). Best results are written bold. The best method with compact representation is written in italic font. Within the methods that have compact representations, $\text{DAIS}_{0}$ achieves the best MAE with respect to the standard deviation in most cases (except for $\text{RBF}_{2},10$ where relative errors are a magnitude smaller in general). In general, $\text{DAIS}_{0}$ performs competitively across all methods. Our theoretical result Theorem 4.1 might explain this, as the Jensen-Shannon divergence (minimized by $\text{DAIS}_{0}$) does not suffer from a mode-seeking behavior but learns a $q_{0}$ that covers more mass of the target distribution. For the mean estimation, IWVI outperforms $\text{DAIS}_{0}$ while $\text{IWVI}_{\text{SIR}}$ performs best. Comparing $\text{DAIS}_{0}$ and DAIS, we find that a compact representation indeed helps in some cases (e.g., mean in $\text{RBF}_{1},d=10,25$). $\text{MSC}_{8\text{c}}$, minimizing the forward KL divergence, performs competitively but slightly worse in comparison to the other methods. We attribute this finding to the instability of the forward KL divergence during optimization. We provide a larger study in Appendix C.
@@ 6 Conclusion
In this work, we investigate the initial distribution of DAIS. We show theoretically that, for many transition steps, it minimizes the Jensen-Shannon divergence (symmetrized KL divergence) to the target distribution. Therefore, by using $q_{0}$ as an approximate posterior (a method that we call $\text{DAIS}_{0}$), we expect it to be less prone to underestimating variances (compared to VI) and easier to optimize in higher-dimensional settings (compared to MSC). Additionally, we argue that $q_{0}$ is an explicit and compact representation of the target distribution, which provides advantages over sampling based methods for various downstream tasks. In experiments on both synthetic and real-world data, we verify that $\text{DAIS}_{0}$ indeed often fulfills our expectations by finding distributions with more accurate variances in higher dimensions (compared to other compact and sampling based methods). While $\text{DAIS}_{0}$ is more expensive than VI at training time, inference with $\text{DAIS}_{0}$ is just as expensive as VI but it is significantly cheaper than DAIS.
@@ Appendix A Proof of the Main Theoretical Result
We assume that $\operatorname{support}(q_{0})\supset\operatorname{support}(f)$, perfect transitions between two annealing distributions and equally spaced $\beta_{k}$, i.e., $\beta_{k}=k/K$. Then, for large $K$ and $N=1$, the gap $\Delta_{\textup{AIS}}^{1,K}\coloneqq\log Z-\mathrm{ELBO}_{\textup{AIS}}^{1,K}(f_{\textup{AIS}},q_{\textup{AIS}})$ is a divergence between the initial distribution $q_{0}$ and the target distribution $f/Z$. Up to corrections of order ${\mathcal{O}}(1/K^{3})$, this divergence is proportional to the Jensen-Shannon divergence (symmetrized KL-divergence),
@@
where $D_{\text{JS}}$ denotes the Jensen-Shannon divergence. This shows that the bound for $N>1$ can be upper-bounded by the Jensen-Shannon divergence but does not give additional insights on, e.g., symmetry. For $N=1$ the inequality is an equality.

### v2

# Differentiable Annealed Importance Sampling Minimizes The Symmetrized Kullback-Leibler Divergence Between Initial and Target Distribution

@@ Abstract
Differentiable annealed importance sampling (DAIS), proposed by Geffner & Domke (2021) and Zhang et al. (2021), allows optimizing over the initial distribution of AIS. In this paper, we show that, in the limit of many transitions, DAIS minimizes the symmetrized Kullback-Leibler divergence between the initial and target distribution. Thus, DAIS can be seen as a form of variational inference (VI) as its initial distribution is a parametric fit to an intractable target distribution. We empirically evaluate the usefulness of the initial distribution as a variational distribution on synthetic and real-world data, observing that it often provides more accurate uncertainty estimates than VI (optimizing the reverse KL divergence), importance weighted VI, and Markovian score climbing (optimizing the forward KL divergence).
@@ 1 Introduction
We show theoretically that, in the limit of many transitions, DAIS minimizes the symmetrized Kullback-Leibler (KL) divergence111 In a previous version, we used the term Jensen-Shannon divergence. To be precise, we have updated our terminology. (Kullback & Leibler, 1951) between its initial and target distribution. The symmetrized KL divergence is the sum of the reverse and the forward KL divergence. While the reverse KL divergence, used in VI, is known to be mode-seeking, the forward KL divergence, used in Markovian Score Climbing (MSC) (Naesseth et al., 2020), is associated with a mass-covering behavior (Murphy, 2023). The symmetrized KL divergence averages between both.
@@
Starting from our theoretical contribution—showing that DAIS minimizes the symmetrized KL divergence between $q_{0}$ and ${f\,/\,Z}$—we provide an empirical analysis of $\text{DAIS}_{0}$ for approximate Bayesian inference. We compare $\text{DAIS}_{0}$, DAIS, VI, importance weighted VI (IWVI), and MSC. We find that $\text{DAIS}_{0}$ often provides more accurate uncertainty estimates than VI, IWVI, and MSC on Gaussian process regression and real-world datasets for logistic regression.
@@ 2 Related Work
In the following, we discuss related work on variational distributions that are augmented by MCMC transitions, the forward and reverse KL divergence, and DAIS. We cover related work on methods to estimate normalization constants in Section 3. Related work on bounds with respect to the symmetrized KL divergence is given in Section 4.
@@
Black box VI (Ranganath et al., 2014) is typically associated with the reverse KL divergence (between the variational distribution and the real posterior distribution). However, also various other divergences have been investigated to obtain an approximation of the posterior distribution (Hernandez-Lobato et al., 2016; Li & Turner, 2016; Dieng et al., 2017; Wan et al., 2020). Most related to $\text{DAIS}_{0}$, Ruiz & Titsias (2019) refine the variational distribution by running MCMC transitions. They minimize a contrastive divergence which, in the limit, converges to the symmetrized KL divergence. Markovian score climbing (MSC) (Naesseth et al., 2020), that we also compare to in Section 5, optimizes the forward KL divergence using unbiased stochastic gradients. MSC samples a Markov chain and uses the samples to follow the score of the variational distribution. The Markov kernel for the MCMC dynamics is based on importance sampling.
@@ 4 Analyzing the Initial Distribution of DAIS
In Theorem 4.1 of this section we present our main theoretical result that DAIS minimizes the symmetrized KL divergence between its initial distribution and its target distribution. Then, we discuss compact representations in terms of their sampling complexity. Finally, we frame finding the initial distribution of DAIS as a form of VI.
4.1 DAIS Minimizes the Symmetrized KL Divergence
Theorem 4.1 below presents our main theoretical contribution showing that DAIS minimizes the symmetrized KL divergence between its initial and target distribution. We see Theorem 4.1 as starting point to motivate $\text{DAIS}_{0}$ and the analysis of the shape of the fitted initial distribution $q_{0}$. The result also holds true for AIS, but it is most relevant for DAIS, which learns parameters of its initial distribution. One can also show Theorem 4.1 by combining results of Grosse et al. (2013) and Crooks (2007) from the physics literature. Brekelmans et al. (2022) state a related bound on the difference between an upper and a lower bound on the evidence (whereas Theorem 4.1 makes an asymptotic statement about the gap between lower bound and evidence).
@@ 4.1 DAIS Minimizes the Symmetrized KL Divergence
Let $\operatorname{support}(q_{0})\supset\operatorname{support}(f)$. We assume that the transitions between consecutive annealing distributions are perfect and that $\beta_{k}$ are equally spaced, i.e., $\beta_{k}=k/K$. Then, for large $K$ and $N=1$, the gap $\Delta_{\mathrm{AIS}}^{1,K}\coloneqq\log Z-\mathrm{ELBO}_{\mathrm{AIS}}^{1,K}(f_{\mathrm{AIS}},q_{\mathrm{AIS}})$ is a divergence between the initial distribution $q_{0}$ and the target distribution $f/Z$. Up to corrections of ${\mathcal{O}}(1\,/\,K^{3})$, the divergence is proportional to the symmetrized KL divergence,
@@
where $D_{\text{JS}}$ denotes the symmetrized KL divergence. This shows that the bound for $N>1$ can be upper-bounded by the symmetrized KL divergence but does not give additional insights on, e.g., symmetry. For $N=1$ the inequality is an equality. See Appendix A for a discussion in greater detail.
@@
In contrast, the forward KL divergence $D_{\mathrm{KL}}({f\,/\,Z}\,\|\,q)$ is mass-covering since the expectation is taken over $f/Z$ and regions in ${\bm{z}}$-space where $q({\bm{z}})<f({\bm{z}})\,/\,Z$ are weighted higher than the regions where $q({\bm{z}})>f({\bm{z}})\,/\,Z$. While the forward KL divergence is less likely to underestimate posterior variances, it often performs poorly in moderate to high dimensions when posterior correlations increase (Dhaka et al., 2021). Section 5 provides empirical evidence that $\text{DAIS}_{0}$ (implicitly minimizing the symmetrized KL divergence) indeed outperforms MSC (minimizing the forward KL divergence).
@@ 4.3 The Initial Distribution of DAIS for Inference
As highlighted in Section 4.2, we do not expect that the initial distribution of DAIS ($\text{DAIS}_{0}$) generally outperforms the DAIS method. However, we highlight that $\text{DAIS}_{0}$ provides various desirable properties due to its compact representation. Additionally, $\text{DAIS}_{0}$ is much more computationally efficient at inference time compared to DAIS. In comparison to VI, $\text{DAIS}_{0}$ is expected be less mode-seeking (Section 4.1), and we pose that the symmetrized KL divergence is less susceptible to problems optimizing the forward KL divergence. Section 5 provides experimental evidence.
@@ 5.1 Bimodal Target Distribution
In Figure 4 we can clearly see that VI and $\text{DAIS}_{0}$ with $K=2$ transitions lead to a mode-seeking distribution. For an increasing $K$ (i.e., $K\in\{4,16\}$), we find that the variational distribution of $\text{DAIS}_{0}$ becomes more mass-covering and increasingly similar to the solution that (numerically) minimizes the symmetrized KL divergence (regularly dashed line). This is consistent with Theorem 4.1. We find that we can unambiguously classify almost all of the learned variational distributions into being mode-seeking (“s”) or mass-covering (“c”) by calculating the distance of the mean of the variational distribution to both modes of the target distribution, and to their mid point (Appendix B provides a histogram of these distances and more details on the classification). Figure 3 shows corresponding classifications as labels “s” and “c” for each combination $(N,K,d)$. We label the few cases where the classification is not perfectly unambiguous with “u” which is short for for “undecidable”.
Comparing (IW)VI and $\text{DAIS}_{0}\ (16)$ in Figure 3, we find that $\text{DAIS}_{0}\ (16)$ typically covers both modes of the target distribution for larger dimensions (across all $N$). We attribute this finding to the symmetrized KL divergence that is implicitly minimized by $\text{DAIS}_{0}$. MSC does not converge to reasonable values (“-”) for $N=1$. With a single chain $\text{MSC}_{1\text{c}}$ outperforms (IW)VI over all $N$ but seems to be less mass-covering than $\text{DAIS}_{0}$. With $8$ parallel chains $\text{MSC}_{8\text{c}}$ sometimes outperforms the other methods in finding a mass-covering distribution. However, we find that the variational distributions of $\text{MSC}_{8\text{c}}$ are often not centered on the mid-point between the two modes which we attribute to the high variance of the method (see Appendix B). This is in accordance with Kim et al. (2022) who report high variance for MSC and Dhaka et al. (2021) who note that the forward KL divergence is difficult to optimize for large dimensions.
@@ 5.2 Gaussian Processes Regression
Table 1 shows mean absolute errors (MAEs) for a larger set of GP models. We compare models with compact distributions (IWVI, $\text{DAIS}_{0}$, $\text{MSC}_{8\text{c}}$) and sampling based approaches (DAIS using $10^{5}$ and $\text{IWVI}_{\text{SIR}}$ using $10^{3}\times 10^{3}$ samples). Best results are written bold. The best method with compact representation is written in italic font. Within the methods that have compact representations, $\text{DAIS}_{0}$ achieves the best MAE with respect to the standard deviation in most cases (except for $\text{RBF}_{2},10$ where relative errors are a magnitude smaller in general). In general, $\text{DAIS}_{0}$ performs competitively across all methods. Our theoretical result Theorem 4.1 might explain this, as the symmetrized KL divergence (minimized by $\text{DAIS}_{0}$) does not suffer from a mode-seeking behavior but learns a $q_{0}$ that covers more mass of the target distribution. For the mean estimation, IWVI outperforms $\text{DAIS}_{0}$ while $\text{IWVI}_{\text{SIR}}$ performs best. Comparing $\text{DAIS}_{0}$ and DAIS, we find that a compact representation indeed helps in some cases (e.g., mean in $\text{RBF}_{1},d=10,25$). $\text{MSC}_{8\text{c}}$, minimizing the forward KL divergence, performs competitively but slightly worse in comparison to the other methods. We attribute this finding to the instability of the forward KL divergence during optimization. We provide a larger study in Appendix C.
@@ 6 Conclusion
In this work, we investigate the initial distribution of DAIS. We show theoretically that, for many transition steps, it minimizes the symmetrized KL divergence to the target distribution. Therefore, by using $q_{0}$ as an approximate posterior (a method that we call $\text{DAIS}_{0}$), we expect it to be less prone to underestimating variances (compared to VI) and easier to optimize in higher-dimensional settings (compared to MSC). Additionally, we argue that $q_{0}$ is an explicit and compact representation of the target distribution, which provides advantages over sampling based methods for various downstream tasks. In experiments on both synthetic and real-world data, we verify that $\text{DAIS}_{0}$ indeed often fulfills our expectations by finding distributions with more accurate variances in higher dimensions (compared to other compact and sampling based methods). While $\text{DAIS}_{0}$ is more expensive than VI at training time, inference with $\text{DAIS}_{0}$ is just as expensive as VI but it is significantly cheaper than DAIS.
@@ Appendix A Proof of the Main Theoretical Result
We assume that $\operatorname{support}(q_{0})\supset\operatorname{support}(f)$, perfect transitions between two annealing distributions and equally spaced $\beta_{k}$, i.e., $\beta_{k}=k/K$. Then, for large $K$ and $N=1$, the gap $\Delta_{\textup{AIS}}^{1,K}\coloneqq\log Z-\mathrm{ELBO}_{\textup{AIS}}^{1,K}(f_{\textup{AIS}},q_{\textup{AIS}})$ is a divergence between the initial distribution $q_{0}$ and the target distribution $f/Z$. Up to corrections of order ${\mathcal{O}}(1/K^{3})$, this divergence is proportional to the symmetrized KL divergence,
@@
where $D_{\text{JS}}$ denotes the symmetrized KL divergence. This shows that the bound for $N>1$ can be upper-bounded by the symmetrized KL divergence but does not give additional insights on, e.g., symmetry. For $N=1$ the inequality is an equality.